OpenStudy (anonymous):

Simplify. (if its already in simplest form, enter given expression) x^2−16/ = (x+4)(x−4)/ a(x+4)−b(x+4) a(x+4)-b(x+4) Please helpe

OpenStudy (anonymous):

thought the answer would be (x+4)(x-4)/a-b (but you cannot simplify the denominator since it has the minus/subtraction symbol correct? if you could help me understand the rule greatly appreciate it )

OpenStudy (anonymous):

(x+4)(x-4)/(a-b)(x+4)=(x-4)/(a-b)

OpenStudy (anonymous):

in denominator (x+4) is common so we can write (x+4)(a-b)

OpenStudy (anonymous):

then (x+4) in numerator is divided by (x+4) in denominator = 1 does not affect our solution